5. The coordinates of the vertex of the curve y = 1 - 2x - 4x2 are:


6. If 2x2 - 12x + 11 ≡ 2(x + b)2 + c, then

                A. b = 6, c = -7               B. b = 3, c = -7          C. b = -3, c = -7               D. b = 3, c = 7


1
Expert's answer
2021-02-24T06:42:12-0500

Ans 5:-

Given curve is-

"y=1-2x-4x^2"


"y=\dfrac{-1}{4}(x^2+\dfrac{x}{2}-\dfrac{1}{4})"


"y=-\dfrac{1}{4}(x^2+\dfrac{x}{2}+\dfrac{1}{16}-\dfrac{1}{16}-\dfrac{1}{4})"



"y=-\dfrac{1}{4}(x+\dfrac{1}{4})^2+\dfrac{5}{64}"


On Comparing the above equation with "y=a(x-h)^2+k"


We get "h=\dfrac{-1}{4},k=\dfrac{5}{64}"


We get the vertex as "(-\dfrac{1}{4},\dfrac{5}{64})"


Ans 6:-

Given, "2x^2 - 12x + 11 = 2(x + b)^2 + c"


"2x^2-12x+11=2(x^2+b^2+2xb)+c\\2x^2-12x+11=2x^2+2b^2+4xb+c"


On comparing coefficient of Both the sides

we get "4b=-12,2b^2+c=11"


"b=\dfrac{-12}{4}=-3"


"\Rightarrow 2b^2+c=11\\\Rightarrow c=11-2(-3)^2\\\Rightarrow c=11-2(9)\\\Rightarrow c=11-18=-7"


Hence The value of "b=-3 \text{ and } c=-7."